# exampledefaultouterlinewidth=0pt,innerlinewidth=0pt, roundcorner=5pt,backgroundcolor=myblue, leftmargin=0.2cm,rightmargin=0.2cm

### v1

@@ 3 Idealized Exact Line Search and Polyak Stepsize
In this section, we consider two classical idealized algorithms – exact line search and Polyak stepsize. In deterministic optimization, these learning rate strategies are chosen so that the function value (exact line search) or distance to optimality (Polyak) produces the largest decrease in function value (resp. distance to optimality) at the next iteration. For stochastic algorithms, we can ask this to hold for the deterministic equivalent to the risk $\mathscr{R}(t)$ (resp. distance to optimality, $\mathscr{D}(t)$) since we know that SGD is close to these deterministic equivalents. Thus, the question is: what choice of learning rate decreases the $\mathscr{R}(t)$ (exact line search) and/or $\mathscr{D}(t)$ (Polyak stepsize)? We will restrict to least squares in this section – see Appendix 10.1 and 10.2 for general functions as well as proofs for least squares. These are idealized algorithms because we can not implement them as they require distributional knowledge of $a$ or $X^{\star}$. Despite this, they provide a basis for more practical algorithms.
@@
The latter yields the optimal fixed learning rate (up to absolute constant factors) for a noiseless target on a least squares problem [34, 42]).
@@
In the context of risk, using (12) and noting that $\mathscr{R}(t)=\tfrac{1}{2d}\sum_{i=1}^{d}\lambda_{i}\mathscr{D}_{i}^{2}(t)$, we can find $\gamma_{t}^{\text{line}}\in\text{arg\,min}\mathop{}\!\mathrm{d}\mathscr{R}{(t)}$; i.e., the greedy learning rate that decreases the risk the most in the next iteration. We call this exact line search. Expressions for the learning rates are given in Table LABEL:table:learning_rate_expressions, (c.f. Appendix 10.1 for general losses). Because these come from ODEs, we can use ODE theory to give exact limiting values for the deterministic equivalent of $\mathfrak{g}_{k}^{\text{line}}$.
@@
For a proof and an explicit formula for $\lim_{t\to\infty}\gamma_{t}^{\text{line}}$, see Section 10.2. Hence, we observe that being greedy for the risk, in a sufficiently anisotropic setting, will badly underperform the Polyak stepsize.
Figure 3: Quantities effecting AdaGrad-Norm learning rate. (left): Effect of noise ($\omega=1.0$) on risk (left axis) and learning rate (right axis). Depicted is learning $\tfrac{\text{learning rate}}{\text{asymptotic}}$ so it approaches $1$. (Center, right): Noiseless least squares $(\omega=0)$. As predicted in Prop. 4.2, $\lim_{t\to\infty}\gamma_{t}$ depends on avg. eig. of $K$ ($\text{Tr}(K)/d$) and $\|X_{0}-X^{\star}\|^{2}$ but not $\kappa=\lambda_{\max}/\lambda_{\min}$. See Appendix 12 for simulation details.
@@ 4 AdaGrad-Norm analysis
In this section, we analyze the behavior of AdaGrad-Norm learning rate in the least squares setting (see Sec. 8 for general strongly convex functions). In the presence of additive noise, the AdaGrad-Norm learning rate decays like $t^{-1/2}$, regardless of the data covariance $K$. In contrast, the model with no noise exhibits a learning rate that depends on the spectrum of $K$. The learning rate is bounded below by a constant when $\lambda_{\min}(K)>0$ is fixed as $d\to\infty$, and we quantify this lower bound. If the limiting spectral measure of $K$ has unbounded density near 0 (e.g. power law spectrum), then the learning rate can approach zero and we quantify the rate of this convergence in the least squares setting as a function of spectral parameters.
@@
An analog of Proposition 4.2 for the strongly convex setting appears in Sec.  8 (see Prop. 8.1). We now consider two cases in which, as $d\to\infty$, there are eigenvalues of $K$ arbitrarily close to 0.
@@
Let $K$ have a spectrum that converges as $d\to\infty$ to the power law measure $\rho(\lambda)=(1-\beta)\lambda^{-\beta}\mathbf{1}_{(0,1)}$, for some $\beta<1$, and suppose that $\mathscr{D}^{2}_{i}(0)\sim\lambda_{i}^{-\delta}$ for $\delta\geq 0$. Then:
@@
This proposition shows non-trivial decay of the learning rate is dictated by the residuals (distance to optimality at initialization) and the spectrum of $K$. We note that $\delta=0$ corresponds to uniform contribution of each mode (e.g. $X_{0}$ normally distributed). As the eigenmodes of the residuals become more localized, the decay of the learning rate is closer to the behaviour in the presence of additive noise. Furthermore, the scaling behaviour of the loss is affected by the structure of the AdaGrad-Norm algorithm (see Fig. 4).
@@
Proofs of the above propositions, in a slightly more general setting, are deferred to Sec. 8.
Outline of the paper.
The remainder of the article is structured as follows: in Section 5 we illustrate two adaptive learning rate algorithms – AdaGrad-Norm and RMSprop-Norm. In Section 6, we prove the main concentration theorem, Theorem 2.1, and an extension to any statistic. The proof of Theorem 2.1 uses that SGD-AL is an approximate solution to the ODEs (7). Indeed, in Section 7, we show this. Moreover, in Section 7.3, we show that AdaGrad-Norm satisfies Assumption 6. In Section 8, we prove all the propositions in Section 4 regarding our analysis of AdaGrad-Norm. We follow this with Section 9 and Section 10 which proves the propositions in Section 3 for Polyak and exact line search, resp. We also provide in these sections the general update rules for Polyak and line search beyond least squares. In Section 11, we illustrate how (noiseless) binary logistic regression problem is in the high-dimensional linear composite class. Finally, in Section 12, we provide our numerical simulation details.
5 SGD adaptive learning rate algorithms and stepsizes
@@
with the deterministic equivalent (see Section 2 and also 7.3) for this learning rate being
@@
Example: RMSprop-Norm.
@@
$\gamma_{t}=\frac{\eta}{\sqrt{b^{2}e^{-\alpha t}+\tfrac{\text{Tr}(K)}{d}\int_{0}^{t}e^{-\alpha(t-s)}I(\mathscr{B}(s))\,\mathop{}\!\mathrm{d}s}},\quad\text{where}\quad\Gamma(t)=\int_{0}^{t}\gamma_{s}\,\mathop{}\!\mathrm{d}s.$ (19)
@@
6 The Dynamical nexus
In this section, we prove the main theorem on concentration of the risk curves and learning rates. We shall set some notation.
In what follows, we use $W=[X|X^{\star}]\in\mathbb{R}^{d\times 2}$. We also use $W^{+}=[W|X_{0}]=[X|X^{\star}|X_{0}]$. Moreover, we use the shorthand $r=\langle a,W\rangle$, and $x=\langle{a,X}\rangle$ so that $f(\langle{a,X}\rangle,\langle{a,X^{\star}}\rangle;\epsilon)=f(\langle{a,W}\rangle;\epsilon)=f(r;\epsilon)$ and we let $B=B(X)=W^{T}KW$ be the covariance matrix of the Gaussian vector $r$. We also write $f^{\prime}$ for the $\partial_{x}f$, i.e., the derivative of $f$ in the first coordinate.
6.1 Discussion of the assumptions on $f$
In this section, we show how the assumptions we put on $h$ and $I$ are almost satisfied for $L$-smooth $f$. We say that $f$ is $L$-smooth if:
@@ 6.1 Discussion of the assumptions on $f$
For the differentiability of $h$, suppose for the moment that $f$ is $C^{2}$ with bounded second derivatives. We shall remove this condition by approximation in what follows. Setting $Q=\sqrt{B}$ the positive semi-definite square root of $B$, we have
@@
This lemma shows that an $L$-smooth function nearly satisfies Assumption 3 and 4 provided that $\|B^{-1}\|_{\text{op}}$ is bounded. Therefore, our concentration result Theorem 6.1 and its corollaries will hold provided we add a stopping time to the theorems/corollaries. In particular, fix $M>0$ and let
@@
Then the concentration of the risk under SGD to a deterministic function, Theorem 6.1, holds with $t$ replaced with $t\wedge\hbar_{M}(B)\wedge\hbar_{M}(\mathscr{B})$. The corollaries of Theorem 6.1 also follow under this added stopping time.
In the next section, we prove this concentration theorem, Theorem 6.1.
6.2 Integro-differential equation for $\mathscr{S}(t,z)$
@@
We first note that there is an actual solution to the integro-differential equation. This solution is the same as the ODEs defined in the introduction (see (7)) and proved in [14, Lemma 4.1].
Lemma 6.2 (Equivalence to coupled ODEs.).
@@
In this section, we will be working with approximate solutions to the integro-differential equation (21) (see below for specifics). To work with these solutions, we introduce some notation. We always work on a fixed contour $\Omega$ surrounding the spectrum of $K$, given by $\Omega\stackrel{{\scriptstyle\text{ def }}}{{=}}\left\{z:|z|=\max\left\{1,2\|K\|_{\text{op}}\right\}\right\}$. We note that this contour is always distance at least $\frac{1}{2}$ from the spectrum of $K$. We define a norm, $\|\cdot\|_{\Omega}$, on a continuous function $A:\mathbb{C}\to\mathbb{R}$ as
@@
Definition 6.1 ($(\varepsilon,M,T)$-approximate solution to the integro-differential equation).
@@
Theorem 6.1 (Concentration of SGD and deterministic function $\mathscr{S}(t,z)$).
@@
By Proposition 7.1, for any $M$ and $T$, we can find a $\tilde{\varepsilon}>0$ such that the function ${S}(W_{td},z)$ is an $(d^{-\tilde{\varepsilon}},M,T)$-approximate solution. (For the deterministic function $\mathscr{S}$, it is an $(0,M,T)$-approximate solution by definition.) We now apply the stability result, [14, Prop. 4.1], to conclude that there exists a $\varepsilon>0$ such that
$\sup_{0\leq t\leq T\wedge\hat{\tau}_{M}}\|\mathscr{S}(t,z)-S(W_{td},z)\|_{{\Omega}}\leq d^{-\varepsilon},\quad w.o.p.$ (31)
The result immediately follows. ∎
Corollary 6.1.
Suppose the assumptions of Theorem 6.1 hold. Let $f$ be an $\alpha$-pseudo-Lipschitz function with $\alpha\leq 1$ and let $q$ be a polynomial. Set
@@
This is basically equivalent to [14, Corollary 4.2]. The only difference is that [14, Corollary 4.2] requires the boundedness of $\mathscr{N}$; however, since our function $f$ is $\alpha$-pseudo-Lipschitz with $\alpha\leq 1$, this boundedness follows from [14, Proposition 1.2], and the rest of the proof is identical to the one in [14]. ∎
Remark 6.1.
The learning rate $\mathfrak{g}_{k}$, technically, is not a function of $W^{T}q(K)W$. However, Assumption 6 ensures that the learning rate concentrates around a function $W^{T}q(K)W$. Therefore, Corollary 6.1 applies to the learning rate.
7 SGD-AL is an approximate solution

### v2

@@ 3 Idealized Exact Line Search and Polyak Stepsize
In this section, we consider two classical idealized algorithms – exact line search and Polyak stepsize. In deterministic optimization, these learning rate strategies are chosen so that the function value (exact line search) or distance to optimality (Polyak) produces the largest decrease in function value (resp. distance to optimality) at the next iteration. For stochastic algorithms, we can ask this to hold for the deterministic equivalent to the risk $\mathscr{R}(t)$ (resp. distance to optimality, $\mathscr{D}(t)$) since we know that SGD is close to these deterministic equivalents. Thus, the question is: what choice of learning rate decreases the $\mathscr{R}(t)$ (exact line search) and/or $\mathscr{D}(t)$ (Polyak stepsize)? We will restrict to least squares in this section – see Appendix F.1 and F.2 for general functions as well as proofs for least squares. These are idealized algorithms because we can not implement them as they require distributional knowledge of $a$ or $X^{\star}$. Despite this, they provide a basis for more practical algorithms.
@@
The latter yields the optimal fixed learning rate (up to absolute constant factors) for a noiseless target on a least squares problem [35, 43]).777 The Polyak stepsize we analyze in this paper differs slightly from the ”classic” stepsize in the literature, that is, $\frac{\mathscr{R}(X_{k})-\mathscr{R}(X^{*})}{||\nabla\mathscr{R}(X_{k})||^{2}}$. Rather than using this form, we skip an approximation step in the derivation [27] and use the exactly optimal form. Both variations of the Polyak stepsize can be analyzed under our assumptions; the choice was admittedly somewhat arbitrary. (Note that in the case of least squares, the two stepsizes coincide.)
@@
In the context of risk, using (8) and noting that $\mathscr{R}(t)=\tfrac{1}{2d}\sum_{i=1}^{d}\lambda_{i}\mathscr{D}_{i}^{2}(t)$, we can find $\gamma_{t}^{\text{line}}\in\text{arg\,min}\mathop{}\!\mathrm{d}\mathscr{R}{(t)}$; i.e., the greedy learning rate that decreases the risk the most in the next iteration. We call this exact line search. Expressions for the learning rates are given in Table LABEL:table:learning_rate_expressions, (c.f. Appendix F.1 for general losses). Because these come from ODEs, we can use ODE theory to give exact limiting values for the deterministic equivalent of $\mathfrak{g}_{k}^{\text{line}}$.
@@
For a proof and explicit formula for $\lim_{t\to\infty}\gamma_{t}^{\text{line}}$, see Section F.2. Hence, being greedy for the risk in a sufficiently anisotropic setting will badly underperform Polyak stepsize (see Fig. 2).
Figure 3: Quantities effecting AdaGrad-Norm learning rate. (left): Effect of noise ($\omega=1.0$) on risk (left axis) and learning rate (right axis). Depicted is learning $\tfrac{\text{learning rate}}{\text{asymptotic}}$ so it approaches $1$. (Center, right): Noiseless least squares $(\omega=0)$. As predicted in Prop. 4.2, $\lim_{t\to\infty}\gamma_{t}$ depends on avg. eig. of $K$ ($\text{Tr}(K)/d$) and $\|X_{0}-X^{\star}\|^{2}$ but not $\kappa=\lambda_{\max}/\lambda_{\min}$. See Appendix H for simulation details.
@@ 4 AdaGrad-Norm analysis
In this section, we analyze the behavior of AdaGrad-Norm learning rate in the least squares setting (see Sec. D for general strongly convex functions). In the presence of additive noise, the AdaGrad-Norm learning rate decays like $t^{-1/2}$, regardless of the data covariance $K$. In contrast, the model with no noise exhibits a learning rate that depends on the spectrum of $K$, as illustrated in Figure 3. The learning rate is bounded below by a constant when $\lambda_{\min}(K)>0$ is fixed as $d\to\infty$, and we quantify this lower bound. If the limiting spectral measure of $K$ has unbounded density near 0 (e.g. power law spectrum), then the learning rate can approach zero and we quantify the rate of this convergence in the least squares setting as a function of spectral parameters.
@@
An analog of Proposition 4.2 for the strongly convex setting appears in Sec.  D (see Prop. D.1). We now consider two cases in which, as $d\to\infty$, there are eigenvalues of $K$ arbitrarily close to 0.
@@
[…]
@@
This proposition shows non-trivial decay of the learning rate is dictated by the residuals (distance to optimality at initialization) and the spectrum of $K$. We note that $\delta=0$ corresponds to uniform contribution of each mode (e.g. $X_{0}$ normally distributed). As the eigenmodes of the residuals become more localized, the decay of the learning rate is closer to the behaviour in the presence of additive noise. Furthermore, the scaling behaviour of the loss is affected by the structure of the AdaGrad-Norm algorithm (see Fig. 4). Lastly, constant stepsize SGD yields $\mathscr{R}(t)\asymp t^{\beta+\delta-2}$, with no transition occurring at $\beta+\delta=1$.
@@
Proofs of the above propositions, in a slightly more general setting, are deferred to Sec. D.
5 Conclusions and Limitations
This work studies stochastic adaptive optimization algorithms when data size and parameter size are large, allowing for nonconvex and nonlinear risk functions, as well as data with general covariance structure. The theory shows a concentration of the risk, the learning rate and other key functions to a deterministic limit, which is described by a set of ODEs. The theory is then used to derive the asymptotic behavior of the AdaGrad-Norm and idealized exact line search on strongly convex and least square problems, revealing the influence of the covariance matrix structure on the optimization. A potential extension of this work would be to study other adaptive algorithms such as D-adaptation, DOG, and RMSprop which are covered by the theory. Studying the asymptotic behavior of the risk and the learning rate may improve our understanding of the performance and scalability of these algorithms on more realistic data. Another important application of the theory would be to analyze the ODEs presented here on nonconvex problems.
The current form of the theory is limited to Gaussian data, though many parts of the proof can be extended easily beyond Gaussian data. The main ODE comparison theorem is also only tuned for analyzing problem setups where the trace of the covariance is on the order of the ambient dimension; when the trace of the covariance is much smaller than ambient dimension, other stepsize scalings of SGD are needed. In addition, the analysis is limited to the streaming stochastic adaptive methods. We conjecture that a similar deterministic equivalent holds also for multi-pass algorithms at least for convex problems. This has already been shown in the least square problem for SGD with a fixed deterministic learning rate [43, 45]. Lastly, numerical simulations on real datasets (e.g., CIFAR-5m) suggests that the predicted risk derived by our theory matches the empirical risk of multipass SGD beyond Gaussian data (see for example Figure 6).
Acknowledgments and Disclosure of Funding
E. Collins-Woodfin was supported by Fonds de recherche du Québec – Nature et technologies (FRQNT) postdoctoral training scholarship and Centre de recherches mathématiques (CRM) Applied math postdoctoral fellowship. Research of B. García Malaxechebarría was in part funded by NSF DMS 2023166 (NSF TRIPODS II). Research by E. Paquette was supported by a Discovery Grant from the Natural Science and Engineering Council (NSERC). C. Paquette is a Canadian Institute for Advanced Research (CIFAR) AI chair, Quebec AI Institute (MILA) and a Sloan Research Fellow in Computer Science (2024). C. Paquette was supported by a Discovery Grant from the Natural Science and Engineering Research Council (NSERC) of Canada, NSERC CREATE grant Interdisciplinary Math and Artificial Intelligence Program (INTER-MATH-AI), Google research grant, and Fonds de recherche du Québec – Nature et technologies (FRQNT) New University Researcher’s Start-Up Program. Additional revenues related to this work: C. Paquette has 20% part-time employment at Google DeepMind.
Supplementary material
Broader Impact Statement. The work presented in this paper is foundational research and it is not tied to any particular application. The set-up is on a simple well-studied high-dimensional linear composites (e.g., least squares, logistic regression, phase retrieval) with synthetic data and solved using known algorithms, e.g., AdaGrad-Norm. We present deterministic dynamics for the training loss and adaptive stepsizes. The results are theoretical and we do not anticipate any direct ethical and societal issues. We believe the results will be used by machine learning practitioners and we encourage them to use it to build a more just, prosperous world.
Appendix A SGD adaptive learning rate algorithms and stepsizes
@@
with the deterministic equivalent (see Section 2 and also C.3) for this learning rate being
@@
Example: RMSprop-Norm
@@
$\gamma_{t}=\frac{\eta}{\sqrt{b^{2}e^{-\alpha t}+\tfrac{\text{Tr}(K)}{d}\int_{0}^{t}e^{-\alpha(t-s)}I(\mathscr{B}(s))\,\mathop{}\!\mathrm{d}s}}.$ (18)
@@
Appendix B The Dynamical nexus
In this section, we prove the main theorem on concentration of the risk curves and learning rates. We shall set some notation. In what follows, we again use $W=[X|X^{\star}]\in\mathbb{R}^{d\times 2}$. We also use $W^{+}=[W|X_{0}]=[X|X^{\star}|X_{0}]$.
We shall also use the shorthand $r=\langle a,W\rangle$, and $x=\langle{a,X}\rangle$ so that $f(\langle{a,X}\rangle,\langle{a,X^{\star}}\rangle;\epsilon)=f(\langle{a,W}\rangle;\epsilon)=f(r;\epsilon)$.
We shall let $B=B(X)=W^{T}KW$ be the covariance matrix of the Gaussian vector $r$. We also write $f^{\prime}$ for the $\partial_{x}f$.
B.1 Discussion of the assumptions on $f$
In this section we show how the assumptions we put on $h$ and $I$ are almost satisfied for $L$-smooth $f$. We say that $f$ is $L$-smooth if:
@@ B.1 Discussion of the assumptions on $f$
For the differentiability of $h$, suppose for the moment that $f$ is $C^{2}$ with bounded second derivatives. [^9] This condition can be removed in a standard way: one creates an $f_{\epsilon}$ which is an approximation to $f$ formed by convolving with an isotropic Gaussian of variance $\epsilon$. This is $C^{2}$ and has bounded second derivatives (as $f$ was smooth). One then takes the limit as $\epsilon\to 0$. Setting $Q=\sqrt{B}$ the positive semi-definite square root of $B$, we have
@@
This lemma shows that an $L$-smooth function nearly satisfies Assumption 3 and 4 provided that $\|B^{-1}\|_{\text{op}}$ is bounded. Therefore, our concentration result Theorem B.1 and its Corollaries will hold provided we add a stopping time. Fix $M>0$ and let
@@
Then the concentration of the risk under SGD to a deterministic function, Theorem B.1, holds with $t$ replaced with $t\wedge\hbar_{M}(B)\wedge\hbar_{M}(\mathscr{B})$. The corollaries of Theorem B.1 also follow under this added stopping time.
In the next section, we prove this concentration theorem, Theorem B.1.
B.2 Integro-differential equation for $\mathscr{S}(t,z)$
@@
We first note that there is an actual solution to the integro-differential equation. This solution is the same as the ODEs defined in the introduction (see (9)) and proved in [15, Lemma 4.1].
Lemma B.2 (Equivalence to coupled ODEs.).
@@
In this section, we will be working with approximate solutions to the integro-differential equation (20) (see below for specifics). For working with these solutions, we introduce some notation. We shall always work on a fixed contour $\Omega$ surrounding the spectrum of $K$, given by $\Omega\stackrel{{\scriptstyle\text{ def }}}{{=}}\left\{z:|z|=\max\left\{1,2\|K\|_{\text{op}}\right\}\right\}$. We note that this contour is always distance at least $\frac{1}{2}$ from the spectrum of $K$. We define a norm, $\|\cdot\|_{\Omega}$, on a continuous function $A:\mathbb{C}\to\mathbb{R}$ as
@@
Definition B.1 ($(\varepsilon,M,T)$-approximate solution to the integro-differential equation).
@@
Theorem B.1 (Concentration of SGD and deterministic function $\mathscr{S}(t,z)$).
@@
By Proposition C.1, for any $M$ and $T$, we can find a $\tilde{\varepsilon}>0$ such that the function ${S}(W_{td},z)$ is an $(d^{-\tilde{\varepsilon}},M,T)$-approximate solution. (For the deterministic function $\mathscr{S}$, it is an $(0,M,T)$-approximate solution by definition.) We now apply the stability result, [15, Prop. 4.1], to conclude that there exists a $\varepsilon>0$ such that
$\sup_{0\leq t\leq T\wedge\hat{\tau}_{M}}\|\mathscr{S}(t,z)-S(W_{td},z)\|_{{\Omega}}\leq d^{-\varepsilon},\quad w.o.p,$ (30)
where $\hat{\tau}_{M}$ is shorthand for $\hat{\tau}_{M}(S(W,\cdot))\wedge\hat{\tau}_{M}(\mathscr{S})$. The result immediately follows. ∎
Corollary B.1.
Suppose the assumptions of Theorem B.1 hold. Let $f$ be an $\alpha$-pseudo-Lipschitz function with $\alpha\leq 1$ and let $q$ be a polynomial. Set
@@
This is basically equivalent to [15, Corollary 4.2]. The only difference is that [15, Corollary 4.2] requires the boundedness of $\mathscr{N}$; however, since our function $f$ is $\alpha$-pseudo-Lipschitz with $\alpha\leq 1$, this boundedness follows from [15, Proposition 1.2], and the rest of the proof is identical to the one in [15]. ∎
Remark B.1.
The learning rate $\mathfrak{g}_{k}$, technically, is not a function of $W^{T}q(K)W$. However, Assumption 6 ensures that the learning rate concentrates around a function $W^{T}q(K)W$. Therefore, Corollary B.1 applies to the learning rate.
Appendix C SGD-AL is an approximate solution